# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ 3.1 Center-Time Frame Interpolation (CTFI)
Center-Time Frame Interpolation (CTFI) as shown in Fig. 8 (a), also known as fixed-time interpolation, is a widely adopted learning paradigm in VFI. Here, models are trained on triplets $(I_{0},I_{\frac{1}{2}},I_{1})$ [178, 6, 179, 50, 180], with $I_{0}$ and $I_{1}$ as inputs and $I_{\frac{1}{2}}$ as the GT center-frame. Owing to the simplicity of supervision and precise GT alignment, this paradigm has been dominant in earlier works [92, 28, 29, 16, 76, 50, 36, 59, 51, 113].
@@
Despite its ease of implementation, CTFI suffers from major limitations in real-world scenarios where intermediate frames are required at arbitrary timestamps. Since models are trained to generate only the center-frame at $t{=}\frac{1}{2}$, they inherently lack temporal flexibility for generating frames at other timestamps. For example, to generate a frame at $t{=}\frac{1}{4}$, the model first synthesizes the center-frame, and then recursively generates $\hat{I}_{\frac{1}{4}}$ conditioned on $(I_{0},\hat{I}_{\frac{1}{2}})$. This recursive strategy is inherently sequential, introducing two key drawbacks [5, 79, 181]. First, it increases computational latency and prevents parallel generation, as each intermediate frame depends on the previously synthesized result. Second, it leads to cumulative errors where artifacts in earlier frames propagate through the inference chain, degrading temporal consistency and overall quality. Additionally, CTFI restricts the temporal upsampling factor to powers of two ($2^{n}$), thereby limiting adaptability in diverse frame-rate conversion scenarios such as real-time video streaming or arbitrary slow-motion synthesis.
3.2 Arbitrary-Time Frame Interpolation (ATFI)
@@
In contrast, Arbitrary-Time Frame Interpolation (ATFI) or multi-frame interpolation as shown in Fig. 8 (b), generalizes the task by enabling interpolation at any arbitrary $t{\in}(0,1)$ between two given frames [5, 30, 7, 76, 75, 62, 35, 79, 59, 53, 54, 45, 114, 161]. This paradigm explicitly receives $t$ as input during training and inference, allowing direct synthesis of frames at specified timestamps and supporting continuous-time interpolation. While some earlier methods [7, 73] perform iterative ATFI in a frame-by-frame fashion, such methods often suffer from temporal jitter due to a lack of continuity modeling. In contrast, temporally-aware models [62, 75] predict multiple intermediate frames in one pass, promoting sequence-level coherence and computational efficiency.
@@
Despite its flexibility, ATFI also presents challenges. Training requires HFR datasets to supervise intermediate frames at diverse timestamps. Additionally, ATFI inherently prone to the velocity ambiguity problem, where multiple plausible motion trajectories can lead to the same intermediate position. This often leads models to average over alternatives, resulting in temporal blur. Furthermore, ATFI must account for non-linear motion such as acceleration or abrupt direction changes, phenomena not easily handled under constant-velocity assumptions. These issues are discussed in depth in Sec. 4.4. Despite these challenges, ATFI remains a versatile and powerful paradigm for real-world applications, offering improved flexibility for slow-motion generation, dynamic frame-rate adaptation, and user-controllable playback.
3.3 Training Strategy
3.3.1 CTFI Training Strategy (CTFI-TS)
CTFI-TS builds training triplets $(I_{0},I_{t},I_{1})$ with $I_{t}$ positioned precisely at the center-point between $I_{0}$ and $I_{1}$. These triplets can be generated by sampling three consecutive frames which are uniformly sampled as shown in Fig. 9 (a). This enables the construction of large-scale training datasets without dense manual annotation. During training, models are supervised exclusively at $t{=}0.5$, and no explicit temporal encoding is involved. At inference, the model is similarly evaluated by predicting center-frames at each stride. While efficient, this strategy inherently limits generalization to other timestamps and requires recursive processing for arbitrary-time synthesis.
@@ 3.3.1 CTFI Training Strategy (CTFI-TS)
Figure 9: Comparison of CTFI-TS and ATFI-TS. (a) CTFI-TS samples exactly three uniformly spaced frames per training example, with only the center-frame used as supervision. (b) ATFI-TS uses $(n{+}1)$ uniformly spaced frames from HFR videos, allowing an intermediate frame at arbitrary timestamp $t\in(0,1)$ to serve as supervision target.
3.3.2 ATFI Training Strategy (ATFI-TS)
@@
ATFI-TS constructs training samples from $(n{+}1)$ consecutive frames, using the first and last as inputs $(I_{0},I_{1})$ and the $(n{-}1)$ intermediate frames as supervision targets for their respective times $t\in(0,1)$. Each $t$ is either provided directly or encoded via temporal embeddings [81, 54, 106]. When HFR videos are available, training data can be flexibly constructed by uniformly sub-sampling frames at a desired interval as shown in Fig. 9 (b). As long as the original frame rate of the video is divisible by the desired interpolation factor, any pair of frames can be selected as inputs, and the frames that lie temporally between them can serve as GT supervision targets. This strategy allows models to learn from a wide distribution of motions and time intervals. Inference is fully parallelizable, frames at any $t\in(0,1)$ can be generated independently, making this approach highly efficient and scalable for real-time and high-frame-rate applications. By explicitly modeling time and enabling continuous supervision, ATFI-TS forms the backbone of modern interpolation frameworks seeking generalizability, temporal coherence, and fine-grained control.
3.4 Loss Functions
@@ 3.4.1 Reconstruction Loss
Reconstruction losses supervise the model to minimize the pixel-wise discrepancy between the predicted intermediate frame $\hat{I}_{t}$ and the GT frame $I_{t}^{\text{GT}}$. These losses are typically applied in the RGB space.
$\mathcal{L}_{1}$ Loss is defined as:
$\mathcal{L}_{1}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{1},$ (18)
@@
which computes the pixel-wise absolute difference between frames.
$\mathcal{L}_{2}$ loss is defined as:
$\mathcal{L}_{2}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{2}^{2},$ (19)
@@
this loss computes the squared error, yielding smoother gradients but often producing overly smoothed outputs, particularly in high-frequency regions or under motion-induced misalignments [106].
Charbonnier Loss [182] is a differentiable variant of the $\mathcal{L}1$ loss:
$\mathcal{L}_{\text{char}}=\rho(I_{t}^{GT}-\hat{I}_{t}),$ (20)
where $\rho(x)=(x^{2}+\epsilon^{2})^{\alpha}$ is the Charbonnier function, with a small constant $\epsilon$ (typically $10^{-3}$) for numerical stability and $\alpha=0.5$. The loss provides smoother gradients than the $\mathcal{L}_{1}$ loss. Owing to its smooth gradient profile and outlier resilience, Charbonnier loss is frequently adopted in VFI for its balanced sensitivity to both sharp detail and robust training stability.
Laplacian loss [183] compares the Laplacian pyramid decompositions of the interpolated and GT frames to supervise frame synthesis across multiple spatial scales:
$\mathcal{L}_{\text{lap}}=\sum_{i=1}^{l}2^{i-1}\left\lVert L^{i}(\hat{I}_{t})-L^{i}((I_{t}^{GT})\right\rVert_{1},$ (21)
@@
where $L^{i}(\cdot)$ is the $i$-th pyramid level. This encourages alignment of both global structure and fine detail, and is often used in conjunction with $\mathcal{L}_{1}$ loss.
Census loss [184], also referred to as ternary loss, evaluates the structural consistency of local image patches under census transformation [185]. It is defined as:
$L_{\text{cen}}=\psi(I_{t}^{GT},\hat{I}_{t}),$ (22)
@@
where $\psi(\cdot,\cdot)$ is a Hamming-like distance function over census-encoded patches. Due to its robustness against illumination and photometric noise, census loss improves particularly effective in unsupervised or self-supervised VFI frameworks.
3.4.2 Perceptual Loss
@@
To enhance perceptual realism, VFI models often incorporate high-level perceptual losses in addition to pixel-wise criteria. A widely adopted formulation computes feature-level distances using a pre-trained VGG network [23]:
$\mathcal{L}_{\text{per}}=\left\lVert\phi(\hat{I}_{t})-\phi(I_{t}^{GT})\right\rVert_{2}^{2},$ (23)
where $\phi$ denotes the feature extractor. This loss promotes structural consistency and encourages synthesis of semantically aligned textures, especially in challenging visual regions.
3.4.3 Adversarial Loss
To further improve realism, adversarial learning frameworks employ a discriminator $D$ trained to distinguish interpolated frames from real ones. The standard GAN objective is:
$\mathcal{L}_{\text{GAN}}=\mathbb{E}_{I_{t}^{\mathrm{GT}}}[\log D(I_{t}^{\mathrm{GT}})]+\mathbb{E}_{\hat{I}_{t}}[\log(1-D(\hat{I}_{t}))].$ (24)
By optimizing this objective jointly with reconstruction losses, the generator learns to produce sharper and more plausible frames. To enforce temporal consistency, recent works also adopt temporal discriminators, which operate on sequences to distinguish coherent dynamics [74, 62].
3.4.4 Flow Loss
@@
Smoothness Loss [16] encourages piecewise smooth flow by penalizing abrupt spatial changes:
$\mathcal{L}_{smooth}=\left\lVert\nabla\mathcal{V}_{0\to 1}\right\rVert_{1}+\left\lVert\nabla\mathcal{V}_{1\to 0}\right\rVert_{1}.$ (25)
Warping Loss [5] measures the reconstruction error after warping one frame to the other using estimated flow:
$\mathcal{L}_{warp}=\|I_{0}-\mathcal{W}(I_{1},\mathcal{V})\|_{1}+\|I_{1}-\mathcal{W}(I_{0},\mathcal{V})\|_{1},$ (26)
where $\mathcal{W}$ denotes the warping operator.
First-order Edge-aware Smoothness Loss [79] is designed to preserve sharp motion discontinuities, this loss attenuates regularization near edges:
$\mathcal{L}_{edge}=\sum_{i=0,1}\exp\left(-e^{2}\sum_{c}|\nabla_{x}I^{0}_{tc}|\right)^{\top}\cdot|\nabla_{x}\mathcal{V}^{0}_{ti}|,$ (27)
where edge strengths are computed via image gradients and used to modulate the smoothness penalty.
4 VFI Challenges
Figure 10: Representative challenges in VFI. (a) Large motion makes it difficult to establish accurate correspondences across frames, especially in cases involving fast-moving objects, deformable structures, or significant camera motion. (b) Occlusion introduces ambiguity, as some regions in the intermediate frame are not visible in either of the input frames, making it unclear what content should be synthesized. (c) Lighting variations, such as shadows, reflections, or changes in illumination, violate brightness constancy assumptions and hinder accurate motion estimation. (d) Non-linear motion refers to changes in motion speed or direction over time, making it difficult to infer intermediate positions.
@@ 4.1 Large Motion
Large motion refers to scenarios where objects undergo substantial displacement between consecutive frames. As shown in Fig. 10 (a), this includes articulated movements (e.g., a person leaning left to right) or abrupt camera motion, which result in wide spatial shifts across the image plane. Such motion is prevalent in real-world videos and presents a fundamental challenge in VFI due to the difficulty of establishing accurate correspondences over long spatial ranges.
@@
To accurately synthesize an intermediate frame, the model must identify where each pixel from the first frame $(I_{0})$ has moved in the following frame $(I_{1})$ which denotes motion or correspondence estimation. When the motion is small, this is relatively straightforward because corresponding pixels remain close. However, large motion induces long-range dependencies that exceed the receptive field of standard networks. Moreover, appearance changes and occlusions further hinder accurate estimation by introducing discontinuities in motion and visibility. To address this, many VFI models adopt a coarse-to-fine hierarchical framework, where large displacements are estimated at low-resolution (LR) feature maps and progressively refined at higher resolutions. RIFE [59] employs multi-scale residual flow refinement, enabling robust alignment under wide motion ranges. FILM [83] leverages a feature pyramid for flow estimation and lightweight synthesis, explicitly targeting fast motion and blur scenarios. Similarly, IFRNet [81] improves motion encoding through a motion-aware feature extractor and an intermediate flow refinement block. In addition to these designs, some models further enhance alignment under large displacements by leveraging bidirectional motion modeling [5, 58, 56, 57] or attention mechanisms [97, 55, 161]. ABME [58] proposes asymmetric bilateral estimation, predicting forward and backward flows independently to improve robustness under occlusion. BiFormer [55] incorporates deformable attention across bidirectional contexts, enabling the model to dynamically attend to semantically relevant but spatially distant regions, an effective strategy for capturing non-local motion patterns.
Despite architectural differences, these models all share a common objective of expanding the receptive field effectively while maintaining spatial precision. To this end, many methods combine multi-scale refinement, attention-based global matching, and motion-aware modules, enabling them to handle wide-range motion more effectively. Such designs have demonstrated strong performance on benchmarks involving extreme motion, most notably X4K1000FPS [79], which provides 4K videos at 1000fps along with dense GT for fine-grained evaluation. Following the introduction of X4K1000FPS, several HR datasets [186, 187] have been proposed to further benchmark performance under high-speed and large-displacement conditions. By providing more realistic and challenging settings, these datasets have enabled better training and evaluation of VFI models in unconstrained environments. As a result, the availability of such benchmarks has accelerated the development of more robust architectures capable of preserving motion detail and fidelity under large displacements.
4.2 Occlusion

### v3

@@ III-A Center-Time Frame Interpolation (CTFI)
Center-Time Frame Interpolation (CTFI) as shown in Fig. 8 (a), also known as fixed-time interpolation, is a widely adopted learning paradigm in VFI. Here, models are trained on triplets $(I_{0},I_{\frac{1}{2}},I_{1})$ [209, 252, 211, 36, 208], with $I_{0}$ and $I_{1}$ as inputs and $I_{\frac{1}{2}}$ as the GT center-frame. Owing to the simplicity of supervision and precise GT alignment, this paradigm has been dominant in early studies [149, 169, 170, 148, 251, 36, 124, 93, 203, 152].
@@
However, CTFI is inherently limited in generating intermediate frames at arbitrary timestamps. Since models are trained exclusively for the center-frame at $t{=}\frac{1}{2}$, they inherently lack temporal flexibility for generating frames at other timestamps. For example, to generate a frame at $t{=}\frac{1}{4}$, the model first synthesizes $\hat{I}_{\frac{1}{2}}$, and then recursively generates $\hat{I}_{\frac{1}{4}}$ conditioned on $(I_{0},\hat{I}_{\frac{1}{2}})$. This sequential process incurs two key drawbacks [107, 204, 174]. First, it increases computational latency and prevents parallel generation, as each intermediate frame depends on the previously synthesized result. Second, it leads to cumulative errors where artifacts in earlier frames propagate through the inference chain, degrading temporal consistency and overall quality. Additionally, CTFI restricts the temporal upsampling factor to powers of two ($2^{n}$), thereby limiting adaptability in diverse frame-rate conversion scenarios such as real-time video streaming or arbitrary slow-motion synthesis.
III-B Arbitrary-Time Frame Interpolation (ATFI)
@@
In contrast, Arbitrary-Time Frame Interpolation (ATFI) or multi-frame interpolation as shown in Fig. 8 (b), generalizes the task to arbitrary timestamps $t\in(0,1)$ between two given frames [107, 167, 6, 251, 187, 34, 177, 204, 93, 270, 133, 113, 287, 231]. This paradigm explicitly receives $t$ as input, enabling direct and continuous-time interpolation without recursion. Earlier methods [6, 145] perform iterative ATFI in a frame-by-frame fashion, often leading to temporal jitter due to a lack of continuity modeling. In contrast, temporally-aware models [34, 187] predict multiple intermediate frames in one pass, promoting temporal coherence and computational efficiency.
@@
While ATFI offers superior flexibility, it introduces new challenges. First, training requires HFR datasets to provide dense supervision at various timestamps. Second, ATFI is inherently susceptible to the velocity ambiguity problem, where multiple motion trajectories can lead to the same intermediate position. This often leads models to average over alternatives, resulting in temporal blur. Third, ATFI must account for non-linear motion such as acceleration or abrupt direction changes, which are difficult to model under constant-velocity assumptions. These challenges are further analyzed in Sec. IV-D. Despite these issues, ATFI remains a versatile and powerful paradigm for real-world applications, offering improved flexibility for slow-motion generation, dynamic frame-rate adaptation, and user-controllable playback.
@@
Figure 9: Comparison of CTFI-TS and ATFI-TS. (a) CTFI-TS samples exactly three uniformly spaced frames per training example, with only the center-frame used as supervision. (b) ATFI-TS uses (n+1)(n{+}1) uniformly spaced frames from HFR videos, allowing an intermediate frame at arbitrary timestamp $t\in(0,1)$ to serve as supervision target.
III-C Training Strategy
III-C1 CTFI Training Strategy (CTFI-TS)
CTFI-TS builds training triplets $(I_{0},I_{t},I_{1})$ with $I_{t}$ positioned at the center-point between $I_{0}$ and $I_{1}$. These triplets can be generated by uniformly sampling three consecutive frames as shown in Fig. 9 (a). This enables the construction of large-scale training datasets without dense manual annotation. During training, models are supervised exclusively at t=0.5t{=}0.5, and no explicit temporal encoding is involved. At inference, the model predicts only the center-frames at each step. While efficient, this strategy inherently lacks the flexibility to synthesize frames at arbitrary timestamps and requires recursive inference.
III-C2 ATFI Training Strategy (ATFI-TS)
@@
ATFI-TS constructs training samples from (n+1)(n{+}1) consecutive frames, using the first and last as inputs $(I_{0},I_{1})$ and the (n1)(n{-}1) intermediate frames as supervision targets for their respective times $t\in(0,1)$. Each $t$ is either provided directly or encoded via temporal embeddings [120, 133, 102]. When HFR videos are available, training data can be flexibly constructed by uniformly sub-sampling frames at a desired interval as shown in Fig. 9 (b). As long as the original frame rate of the video is divisible by the desired interpolation factor, any pair of frames can be selected as inputs, and the frames that lie temporally between them can serve as GT supervision targets. This strategy allows models to learn from a wide distribution of motions and time intervals. Unlike CTFI-TS, inference in ATFI-TS is fully parallelizable, frames at any $t\in(0,1)$ can be generated independently. By explicitly modeling time and enabling continuous supervision, ATFI-TS forms the backbone of modern interpolation frameworks seeking generalizability, temporal coherence, and fine-grained control.
III-D Loss Functions
@@ III-D1 Reconstruction Loss
Reconstruction losses supervise the model to minimize the pixel-wise discrepancy between the predicted frame $\hat{I}_{t}$ and the GT frame $I_{t}^{\text{GT}}$. These losses are typically applied in the RGB space.
@@
  • $\mathcal{L}_{1}$ Loss computes the pixel-wise absolute difference between frames, defined as: $\mathcal{L}_{1}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{1}$
@@
  • $\mathcal{L}_{2}$ loss computes the squared error, defined as: $\mathcal{L}_{2}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{2}^{2}$, yielding smoother gradients but often producing overly smoothed outputs, particularly in high-frequency regions or under motion-induced misalignments [102].
  • Charbonnier Loss [21] is a differentiable variant of the $\mathcal{L}1$ loss, defined as: $\mathcal{L}_{\text{char}}=\rho(I_{t}^{GT}-\hat{I}_{t})$ where $\rho(x)=(x^{2}+\epsilon^{2})^{\alpha}$ is the Charbonnier function, with a small constant $\epsilon$ (typically $10^{-3}$) and $\alpha=0.5$. The loss provides smoother gradients than the $\mathcal{L}_{1}$ loss. Owing to its smooth gradients and robustness to outliers, it is widely used in VFI for stable optimization and accurate reconstruction.
@@
  • Laplacian loss [16] compares the Laplacian pyramid decompositions of the interpolated and GT frames to supervise frame synthesis across multiple spatial scales: $\mathcal{L}_{\text{lap}}=\sum_{i=1}^{l}2^{i-1}\left\lVert L^{i}(\hat{I}_{t})-L^{i}((I_{t}^{GT})\right\rVert_{1}$ where $L^{i}(\cdot)$ denotes the $i$-th level of the Laplacian pyramid. This encourages alignment of both global structure and fine detail, and is often used in conjunction with $\mathcal{L}_{1}$ loss.
@@
  • Census loss [155], also referred to as ternary loss, evaluates the structural consistency of local image patches under census transformation [264]. It is defined as: $L_{\text{cen}}=\psi(I_{t}^{GT},\hat{I}_{t})$ where $\psi(\cdot,\cdot)$ is a Hamming-like distance function over census-encoded patches. Due to its robustness against illumination and photometric noise, it is particularly effective in unsupervised or self-supervised VFI frameworks.
III-D2 Perceptual Loss
@@
To enhance perceptual realism, VFI models often incorporate high-level perceptual losses in addition to pixel-wise criteria. A widely adopted formulation computes feature-level distances using a pre-trained VGG network [207]: $\mathcal{L}_{\text{per}}=\left\lVert\phi(\hat{I}_{t})-\phi(I_{t}^{GT})\right\rVert_{2}^{2}$ where $\phi$ denotes the feature extractor. This loss promotes structural consistency and encourages synthesis of semantically aligned textures, especially in challenging visual regions.
III-D3 Adversarial Loss
Adversarial loss enhances realism by training a discriminator $D$ to distinguish interpolated frames from GT. The typical GAN objective is: $\mathcal{L}_{\text{GAN}}=\mathbb{E}_{I_{t}^{\mathrm{GT}}}[\log D(I_{t}^{\mathrm{GT}})]+\mathbb{E}_{\hat{I}_{t}}[\log(1-D(\hat{I}_{t}))]$
Figure 10: Representative challenges in VFI. (a) A person bends between $I_{0}$ and $I_{1}$, creating large articulated motion that makes it difficult to establish accurate correspondences. (b) A cyclist is partially hidden by a passing car, illustrating occlusion where intermediate content is not directly visible in either input frame. (c) Illumination changes from shadowed to fully lit, showing that lighting variation breaks brightness constancy and degrades motion estimation. (d) A beach ball moves along a curved trajectory over wavy water, exemplifying non-linear motion and dynamic textures in which both object paths and stochastic wave patterns cannot be well described by simple linear-flow assumptions.
III-D4 Flow Loss
@@
  • Smoothness Loss [148] encourages piecewise smooth flow by penalizing abrupt spatial changes: $\mathcal{L}_{smooth}=\left\lVert\nabla\mathcal{V}_{0\to 1}\right\rVert_{1}+\left\lVert\nabla\mathcal{V}_{1\to 0}\right\rVert_{1}$
  • Warping Loss [107] measures the reconstruction error after warping one frame to the other using estimated flow: $\mathcal{L}_{warp}=\|I_{0}-\mathcal{W}(I_{1},\mathcal{V})\|_{1}+\|I_{1}-\mathcal{W}(I_{0},\mathcal{V})\|_{1}$ where $\mathcal{W}$ denotes the warping operator.
  • First-order Edge-aware Smoothness Loss [204] is designed to preserve sharp motion discontinuities, this loss attenuates regularization near edges: $\mathcal{L}_{edge}=\sum_{i=0,1}\exp\left(-e^{2}\sum_{c}|\nabla_{x}I^{0}_{tc}|\right)^{\top}\cdot|\nabla_{x}\mathcal{V}^{0}_{ti}|$ where edge strengths are computed via image gradients and used to modulate the smoothness penalty.
IV VFI Challenges
Figure 11: Visual comparison of VFI results under representative challenges. Rows correspond to Vimeo90K (occlusion-dominated), Adobe240fps (moderate motion and blur), and X-TEST (large motion). The first column (“Overlap”) shows the two input frames with a blue crop region; the following columns show the ground truth and the outputs of FLAVR (kernel-based), GIMM-VFI (flow-based), and FCVG (DM-based). Yellow, purple, and green arrows denote motion blur, ghosting artifacts, and severely distorted structures, respectively. (*,*) indicate the $50^{\text{th}}$-percentile occlusion and optical flow magnitudes [204].
@@ IV-A Large Motion
Large motion refers to scenarios where objects undergo substantial displacement between consecutive frames, as shown in Fig. 10 (a). This includes articulated movements (e.g., a person leaning left to right), abrupt camera motion, or rapid object translations, all of which cause wide spatial shifts on the image plane. Such motion is common in real-world videos and presents a fundamental challenge in VFI due to the difficulty of establishing accurate correspondences over long spatial ranges.
To accurately synthesize an intermediate frame, the model must identify where each pixel from the first frame $(I_{0})$ has moved in the following frame $(I_{1})$ which denotes motion or correspondence estimation. When the motion is small, this is relatively straightforward because corresponding pixels remain close. However, large motion induces long-range dependencies that exceed the receptive field of standard networks. Moreover, appearance changes and occlusions further hinder accurate estimation by introducing discontinuities in motion and visibility.
@@
To address this, many VFI models adopt a coarse-to-fine hierarchical framework, estimating large displacements at low-resolution (LR) feature maps and progressively refining them at higher resolutions. RIFE [93] employs multi-scale residual flow refinement for robust alignment over wide motion ranges, while FILM [188] leverages a feature pyramid to improve flow estimation in fast motion and blur scenarios. IFRNet [120] enhances motion encoding with a motion-aware feature extractor and an intermediate flow refinement block. In addition to these designs, bidirectional motion modeling [107, 178, 266, 139] and attention mechanisms [46, 176, 231] further improve alignment under extreme motion. ABME [178] proposes asymmetric bilateral estimation, predicting forward and backward flows independently to improve robustness under occlusion. BiFormer [176] incorporates deformable attention across bidirectional contexts, enabling the model to dynamically attend to semantically relevant but spatially distant regions, an effective strategy for capturing non-local motion patterns.
These models all share a common objective of expanding the receptive field effectively while maintaining spatial precision. Combining multi-scale refinement, global attention-based matching, and motion-aware modules has proven especially effective in handling large motion. Their advantages are evident on challenging benchmarks such as X4K1000FPS [204], which offers 4K videos at 1000fps with dense GT for precise evaluation. Following this, several HR datasets [165, 210, 84] have been proposed to further benchmark performance under high-speed and large-displacement conditions. By providing more realistic and challenging settings, these datasets enable better training and evaluation of VFI models in unconstrained environments. As a result, the availability of such benchmarks has accelerated the development of more robust architectures capable of preserving motion detail and fidelity under large displacements. To concretize these behaviors, Fig. 11 (third row) compares a kernel-based model (FLAVR [113]), a flow-based model (GIMM-VFI [73]), and a diffusion-based model (FCVG [240]) on an X-TEST [204] example with extreme motion. Here the kernel-based FLAVR exhibits pronounced motion trails and over-smoothed structures around the turbine hub and blade (yellow arrows), reflecting its limited effective motion range. The flow-based GIMM-VFI substantially reduces blur but still shows noticeable geometric distortions near the blade tip and hub (red arrows), indicating sensitivity to flow errors under very long-range displacements. The DM-based FCVG, in contrast, preserves sharp and coherent blade and hub structures (green arrows) while occasionally deviating slightly from the exact GT contour, as it prioritizes perceptual plausibility over strict pixel-wise alignment. Together with the middle row from Adobe240fps, where motion and occlusion lie between these two regimes and FLAVR tends to show stronger blur around moving objects whereas GIMM-VFI and FCVG maintain relatively clearer object boundaries, these qualitative comparisons indicate that explicit flow modeling and generative refinement both provide advantages over purely kernel-based alignment when large displacements dominate, while still exhibiting characteristic trade-offs between robustness and geometric fidelity.
IV-B Occlusion